The average time spent in a week by grade 11 students browsing their favorite social networking sites is 12 hours with a standard deviation of 3 hours. If a random sampling of 40 students is selected, find the mean, variance, and standard error of the distribution of all sample means that can be derived from the sample.


1
Expert's answer
2022-03-16T04:53:10-0400

Mean of the sampling means is equal to the population mean, so a = 12 hours

Variance: "S^2={\frac V n}={\frac {3^2} {40}}={\frac 9 {40}}"

Standard deviation: "S=\sqrt{{\frac 9 {40}}}={\frac {3\sqrt{10}} {20}}"

standard error: "SE={\frac S {\sqrt {40}}}={\frac 3 {40}}"